OpenStudy (anonymous):

what is the solution of this equation 1w=4 do i divide or multiply

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you do nothing

OpenStudy (misty1212):

\[1w=w=4\]

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

This does not mean that you will do nothing in all the problems you are left with. :P

OpenStudy (anonymous):

but she was right because 1 divide 4 is 4

OpenStudy (anonymous):

Really, 1w=4 means w=4 so 1x4=4 or 4=1=4 so thats what makes 1w=4 thats like saying 1(4) = 4 because you get that say result because it's 1.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

sorry i meant / (divide) 4/1 = 4

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

1w means multiplying 1 with w.. When we multiply 1 with some variable, we get same variable..

OpenStudy (anonymous):

You have to right the equation 1w w = ? 1w = 4 w= 4 or sum like that you get it now?

OpenStudy (anonymous):

\[1w = 1 \times w = w \; \; \; \; itself\]

OpenStudy (anonymous):

Exactly water, thats what you do because 1(4) = 4

OpenStudy (anonymous):

This is hard to explain because im only in 6th grade D:

OpenStudy (anonymous):

I mean................

OpenStudy (anonymous):

\[1w = 4 \\ 1 \times w = 4 \\ w = 4\]

OpenStudy (anonymous):

right